A 11.2 L balloon is filled with H2(g) at 55.12 oC producing a pressure of 1.21 atm. Assuming that the balloon can expand or contract over a sufficient range of volume, what is its volume in L when the temperature is 126.7 oC and the pressure is 1,810 torr?

---please show 3 decimal points
Answer #1

Given

P 1 = 1.21 atm , V 1 = 11.2 L , T 1 = 55.12 + 273 = 328.12 K

P 2 =1810 torr = 2.382 atm , V 2 = ? , T 2 = 126.7 + 273 = 399.7 K

In given case, amount of Hydrogen gas remains constant.

Hence, we can write n 1 = n 2

When n 1 = n 2 We have equation ( P 1 V 1 / T 1 ) = ( P 2 V 2 / T 2 )

On rearranging above equation ,we get

V 2 = V 1 \times ( P 1 / P 2 ) \times ( T 2 / T 1 )

Substituting given values in above equation ,we get

V 2 = 11.2 L \times ( 1.21 atm / 2.382 atm ) \times ( 399.7 K / 328.12 K )

V 2 = 6.93047 L

ANSWER : 6.930 L